OpenStudy (anonymous):

Consider the function f(x)=x^2e^(9x). f(x) has two inflection points at x=C & x=D with C≤D. Find these points.

OpenStudy (anonymous):

solve f"(x) = 0 f'(x) = 2x*e^(9x) + 9e^(9x)*x^2 = e^(x)*x*(2 + 9x) f"(x) = (e^x*x + e^x)(2 + 9x) + 9e^x*x = e^x[(x + 1)(2 + 9x) + 9x] = 0 so (x + 1)(9x + 2) + 9x = 0 so 9x^2 + 20x + 2 = 0

OpenStudy (anonymous):

uhhh...oops...made an error

OpenStudy (anonymous):

f'(x) = e^(9x)*x*(2 + 9x) so f"(x) = (9*e^(9x)*x + e^(9x))(9x + 2) + 9xe^(9x) = e^(9x)[(9x + 1)(9x + 2) + 9x] = 0 (9x + 1)(9x + 2) + 9x = 0 so 81x^2 + 36x + 2 = 0

OpenStudy (anonymous):

yea i got the same answer x=-2/9 and x=i had gotten -1 too bu ti see where i went wrong thanks for the fix

OpenStudy (anonymous):

solutions: - sqrt(2)/9 - 2/9 sqrt(2)/9 - 2/9